# Learning topological states from randomized measurements using variational tensor network tomography

### v3

@@ III.2 Learning a Rydberg quantum spin liquid
where $\Omega$ represents the Rabi frequency, $\Delta$ is the detuning of the laser drive, and $V_{\ell,\,\ell^{\prime}}$ is the van der Waals interaction potential. The strong Rydberg-Rydberg interactions prevent neighboring atoms lying within a “blockade radius”, $R_{b}$, from being simultaneously excited to the Rydberg state, thereby engendering strong quantum correlations. For an appropriate choice of the blockade radius, configurations of Rydberg atoms on the ruby lattice can be mapped to a set of “dimers” on the kagome lattice [Fig. 3(a)]; such quantum dimer models have long been known to host topological quantum spin liquids [49, 50]. In the Rydberg system, depending on the detuning parameter $\Delta/\Omega$, the ground state (for this specifically chosen $R_{b}$) can be either a trivial disordered phase or a topologically ordered $\mathbb{Z}_{2}$ quantum spin liquid phase [51, 52, 7, 53]; see Fig. 3(a). In the computational basis, these two phases are difficult to distinguish as they both lack any symmetry-breaking order. Here, we select two parameters in the phase diagram belonging to the spin-liquid ($\Delta/\Omega=1.7$) or the disordered ($\Delta/\Omega=0.5$) phase and perform DMRG simulations to find these ground states (see Appendix F for more details).
@@
Figure 4: Prediction of spin-liquid properties. (a) A quantum spin liquid (QSL) state can be viewed as a superpostions of dimers in the computational basis. (b) Definition of measurable diagonostic quantities such as the $Z$ string operators. Top panel: A dashed line across a bond is equivalent to applying the Pauli $Z$ operator to the qubit on the same bond. The $Z$ logical string $P_{\mathrm{Logical}}$ applies the $Z$ string operator across the system. Bottom panel: An open string diagonal FM operator $P_{\mathrm{FM}}$, renormalized by the expectation value of a closed-loop operator; e.g., the semicircle corresponds to the product of $6$ Pauli $Z$ operators. (c) The expectation values of the $P_{\mathrm{FM}}$ (red) and $P_{\mathrm{Logical}}$ (blue) operators for different numbers of samples, $N$, converge to their exact values (lines). (d) Predictions (crosses) of the entanglement entropy for different sizes of the bipartition and (e) the Schmidt values (cut at site $i=21$) from MPS ansätze trained with $N$ samples converge to the exact values (open circles).
@@
Importantly, for this system, we do not have prior knowledge of which observables are important to regularize, unlike for the surface code studied above. Hence, analogously to MPS tomography, we consider estimations of all estimable Pauli strings that are supported on a local subsystem (taken to be 6 qubits here) and use them to regularize our MLE loss function. This subsystem should be chosen such that it captures the important correlations of the state and that the Paulis supported on this subsystem can be estimated accurately using randomized measurements  [^1] Note that we could also consider regularization from global measurements, but then we would only be able to estimate $X^{\otimes k}$ or $Z^{\otimes k}$.. With regard to the task of learning the states of $n=48$ qubits, we now show that our protocol (see Appendix E) is capable of learning both regimes in the phase diagram—i.e., the spin liquid as well as the disordered state—which cannot be distinguished by a local order parameter due to the topological nature of the former.
@@ IV Discussion and outlook
In this work, we combine randomized measurements with tensor networks to perform quantum state tomography via MLE, with optional regularization using classical shadows. While the randomized measurements toolbox is useful for estimating many physical properties, tensor networks are known to efficiently represent physically relevant quantum states. Here, we have considered a more restricted measurement setting, random-$XZ$ shadows, that is well motivated by both its tomographic completeness on real and pure states and its feasibility in near-term experiments. Using these $XZ$ measurements, we can regularize our MLE loss function with observables estimated via the classical shadow framework. We showcase the performance of our protocol by learning the topological ground states of the surface code Hamiltonian as well as the Rydberg Hamiltonian on the ruby lattice. For both of these states, our protocol with random-$XZ$ measurements can accurately reconstruct the target state up to a fidelity of over $0.95$ with under $10^{5}$ samples. In addition, we test our shadow regularization technique on these two applications: we find that requiring the stabilizer expectations values to agree with estimations from classical shadows further improves the fidelity for the surface code, whereas for the ruby-lattice Rydberg spin liquid, this only reduces the number of outliers observed during the optimization process.
@@
Using higher-dimensional TNs combined with randomized measurements would, in principle, allow us to learn any states of interest. In the present work, we have numerically benchmarked our protocol using MPS as our model prior, given its advantages of exact contractions and sampling. An immediate extension of our work would be to generalize this model’s architecture to represent $2$D area-law entangled states, such as string bond states [57], projected entangled pair states (PEPS) [58]. We leave the numerical demonstrations of our tomographic method for such tensor-network states to future work.
@@
It constitutes an interesting perspective to extend these methods to the study of fermionic quantum states, which are the focus of attention in the study of correlated electron systems. Our analysis could be extended to quantum gas microscope measurements of optical lattice Fermi-Hubbard systems [59] which yield snapshots with the charge and spin on each site. While in this setting, quench dynamics generated by noninteracting Hamiltonians are presumably most conceivable [60], it is possible that the spin degrees of freedom can be measured using a similar unitary ensemble such as our random-$XZ$ measurements. The variational optimization can be carried out using fermionic neural networks [61] or auxiliary wavefunctions [62, 63], replacing the MPS we have employed in the present work. To enhance the optimization process using neural-network ansätze, it would be especially interesting to extend recent results [42] on locally-randomized single-qubit measurements to such fermionic systems.
@@
While our method can be generalized to quantum states realized in a wide class of quantum processors, more specifically, for Rydberg quantum simulators, it can be easily extended to learning logically encoded quantum states. In particular, our random-$XZ$ measurements are well suited for near-term encodings, such as the color codes recently implemented in neutral atom arrays [31]. In such a Rydberg array, every logical qubit, encoded, for example, in a Steane code, can be measured in either the $X$ or the $Z$ basis, allowing for both error detection via post-selection and the ability to perform high-fidelity random-$XZ$ measurements. Even if current experiments do not aim to perform full logical quantum state tomography, our characterization and derivation of the tools required for random-$XZ$ classical shadows can be immediately applied to these logical encodings.
@@ V Code and data availability
The numerical code used for producing results and generating the figures in this paper is available at: https://github.com/teng10/tn-shadow-qst.
@@ Acknowledgements.
We thank Jacob Barandes, Christian Bertoni, Jonathan Conrad, Zoe Holmes, Robert Huang, Dmitrii Kochkov, Nathan Leitao, Alexander Nietner, Jakob Unfried, Guifre Vidal, and Manuel Rudolph for useful discussions. We thank Hong-Ye Hu and Leandro Aolita for feedback on the manuscript. Y.T. is grateful to Dmitrii Kochkov for his inputs on the initial prototype of the numerical code base and systematic code reviews. Y.T. also thanks the groups of Zoe Holmes and Jens Eisert for their hospitality during her visits, when part of this work was completed. R.S. is supported by the Princeton Quantum Initiative Fellowship. K.V.K. acknowledges support from the Fannie and John Hertz Foundation and the National Defense Science and Engineering Graduate (NDSEG) Fellowship. The Berlin team acknowledges support from the BMBF (MuniQCAtoms), the DFG (CRC 183, FOR 2724), the Munich Quantum Valley, and the ERC (DebuQC). Y.T. and S.S. have been supported by the U.S. Department of Energy under Grant DE-SC0019030.
@@
This appendix aims to provide a comparison of our method to existing quantum state tomography techniques. We highlight the unique challenges and opportunities in characterizing complex quantum systems, particularly two-dimensional topological states. In addition to the references summarized below, we direct interested readers to Ref. [66] for a comprehensive review of quantum state tomography. Finally, having discussed advantages of our method for learning quasi-$2d$ quantum states with topological order, we summarize the limitations faced by our method discussed in the main text.
@@
    Matrix product state (MPS) tomography [18, 19, 67, 20] is an effective method for reconstructing injective one-dimensional quantum states with short-range correlations associated with a local parent Hamiltonian, by finding the best MPS consistent with local reduced density matrices from randomized measurements. In principle, these reduced density matrices can be estimated from Pauli measurements using the techniques of classical shadows [68]. Its application to $2d$ topological states, however, is challenging: first of all, the size of the reduced density matrices required for uniquely determining a quantum state in $2d$ is unclear, and secondly, due to nontrivial ground-state degeneracies of topological states, they cannot be uniquely determined by local operators and thus, cannot be efficiently determined with this method.
@@
    Variational MPS tomography [26, 27, 28] extends MPS tomography to one-dimensional states without the injectivity constraint. It uses a variational approach to optimize the MPS representation through maximum likelihood estimation (MLE). Prior work also considered using matrix product operator (MPO) for quantum process tomography [29]. Our work extends this protocol to $2d$ states using a “snake path”, commonly adopted in cylindrical DMRG calculations for $2d$ systems. We further enhance our method by incorporating regularization techniques inspired by classical shadow tomography, combining the benefits of both approaches to better characterize two-dimensional quantum systems. We refer the reader to Sec. II for the details of our protocol.
@@
    Neural network quantum state tomography (NNQST) [23] utilizes neural networks as variational ansätze within the MLE framework to model quantum states. Although particularly successful for states with fully positive amplitudes (for example, ground states of frustration-free stoquastic Hamiltonians [32]), this method requires careful consideration when the target states have nontrivial signs and phases. Using measurement outcomes from different bases is therefore essential—a requirement directly addressed in MPS-based approaches, which can inherently relate measurements post local unitary transformations to learn relative phases. In this context, it may be useful to explore a most recently proposed architecture that can accurately represent quantum spin liquids by incorporating physical symmetries [69].
@@
    Generative modeling for density matrix reconstruction [24] extends the use of neural networks to mixed-state tomography. This approach uses single-qubit positive operator-valued measure (POVM) measurements for information-complete reconstructions. While generative models can theoretically model measurement probability distributions across an exponentially large outcome space, practical challenges arise in ensuring convergence to the accurate density matrix, particularly given the sparse nature of such distributions in large systems. It is unclear how effective this approach is [25] for learning entangled quantum states such as the topological ordered states considered in this work.
@@ A.1 Potential advantages and limitations
Our variational tensor network tomography, aimed at diagnostics and recovery of quantum states, offers potential for extracting various properties such as the Von Neumann entanglement entropy, difficult to learn using other techniques. This method combines maximum likelihood estimation of the quantum state with randomized measurements and is detailed in previous works such as [26, 28, 29]. We demonstrate this approach using simplified $XZ$ measurements, which proves sufficient for the two-dimensional topological quantum states considered, with complex long-range entanglement. Such quantum states are represented by non-injective MPS, and can not be learned via MPS tomography [18, 19].
One significant limitation is that the target state must admit an efficient tensor network representation. In numerical experiments, we use an MPS ansatz, recognizing that while it effectively approximates one-dimensional states, it may not capture the full complexity of two-dimensional systems. This works sets the stage for future exploration of more sophisticated tensor networks. We also discuss the adaptation of statistical properties of MLE for quantum tomography, outlining both its potential and the challenges posed by the gauge freedom in MPS representations.
Another limitation is that real laboratory states can be mixed, which requires using generalizations of tensor networks such as locally purified matrix product operators [79]. Additionally, one generally do not have access to the ground truth fidelity with the lab state. This necessitates methods like cross-validation as discussed in Appendix G or efficient fidelity certifications [42] to validate the trained model against empirical data.
@@
Our variational algorithm in the framework of maximum likelihood estimation (MLE) aims to find the MPS representation $|\psi({A})\rangle$ that is most consistent with measurements of the target state vector $|\phi\rangle$. Intuitively, we expect that the accuracy of the optimal MLE solution improves with more samples. The goal of this appendix is to analyze the accuracy of such a solution as a function of the number of samples $N$. More precisely, we provide a probabilistic upper bound of the infidelity $1-F$ by deriving a concentration inequality using the statistical properties of MLE. First, we summarize our main finding of the probabilistic bound in the theorem below:
@@
Now, a parameter shifted by a multiples of $2\pi$ still represents the same distribution $\mathcal{P}_{\theta_{0}}=\mathcal{P}_{\theta_{0}+2\pi}$. Therefore, $\hat{\theta}^{(N)}$ need not converge to $\theta_{0}$. In physics, the phenomenon where multiple parameters describe the same function is known as gauge redundancy. In this simple scenario, we can explicitly account for the periodicity in the optimal estimator, or equivalently impose “gauge fixing” conditions to select a unique representative, by restricting the parameter domain to $[0,2\pi)$. However, for distributions parametrized by MPS, it is challenging in practice to fix the gauge degrees of freedom consistently. Specifically, transforming MPS to canonical forms using singular value decomposition only partially fixes the gauge. Additional unitary-gauge degrees of freedom remain in such a decomposition. Moreover, degeneracies in the singular values introduce redundancy within the subspace spanned by degenerate eigenvectors.

### v4

@@ III.2 Learning a Rydberg quantum spin liquid
where $\Omega$ represents the Rabi frequency, $\Delta$ is the detuning of the laser drive, and $V_{\ell,\,\ell^{\prime}}$ is the van der Waals interaction potential between atoms at sites $\ell$ and $\ell^{\prime}$. The strong Rydberg-Rydberg interactions prevent neighboring atoms lying within a “blockade radius”, $R_{b}$, from being simultaneously excited to the Rydberg state, thereby engendering strong quantum correlations. For an appropriate choice of the blockade radius, configurations of Rydberg atoms on the ruby lattice can be mapped to a set of “dimers” on the kagome lattice [Fig. 3(a)]; such quantum dimer models have long been known to host topological quantum spin liquids [49, 50]. In the Rydberg system, depending on the detuning parameter $\Delta/\Omega$, the ground state (for this specifically chosen $R_{b}$) can be either a trivial disordered phase or a topologically ordered $\mathbb{Z}_{2}$ quantum spin liquid phase [51, 52, 7, 53]; see Fig. 3(a). In the computational basis, these two phases are difficult to distinguish as they both lack any symmetry-breaking order. Here, we select two parameters in the phase diagram belonging to the spin-liquid ($\Delta/\Omega=1.7$) or the disordered ($\Delta/\Omega=0.5$) phase and perform DMRG simulations to find these ground states (see Appendix F for more details).
@@
Figure 4: Prediction of spin-liquid properties. (a) A quantum spin liquid (QSL) state can be viewed as a superpositions of dimers in the computational basis. (b) Definition of measurable diagnostic quantities such as the $Z$ string operators. Top panel: A dashed line across a bond is equivalent to applying the Pauli $Z$ operator to the qubit on the same bond. The $Z$ logical string $P_{\mathrm{Logical}}$ applies the $Z$ string operator across the system. Bottom panel: An open-string diagonal Fredenhagen-Marcu operator $P_{\mathrm{FM}}$, renormalized by the expectation value of a closed-loop operator; e.g., the semicircle corresponds to the product of $6$ Pauli $Z$ operators. (c) The expectation values of the $P_{\mathrm{FM}}$ (red) and $P_{\mathrm{Logical}}$ (blue) operators for different numbers of samples, NN, converge to their exact values (lines). (d) Predictions (crosses) of the entanglement entropy for different sizes of the bipartition and (e) the Schmidt values (cut at site $i=21$) from MPS ansätze trained with NN samples converge to the exact values (open circles).
@@
Importantly, for this system, we do not have prior knowledge of which observables are important to regularize, unlike for the surface code studied above. Hence, analogously to MPS tomography, we consider estimations of all estimable Pauli strings that are supported on a local subsystem (taken to be 6 qubits here) and use them to regularize our MLE loss function. This subsystem should be chosen such that it captures the important correlations of the state and that the Paulis supported on this subsystem can be estimated accurately using randomized measurements  [^1] Note that we could also consider regularization from global measurements, but then, we would only be able to estimate $X^{\otimes k}$ or $Z^{\otimes k}$.. With regard to the task of learning the states of $n=48$ qubits, we now show that our protocol (see Appendix E) is capable of learning both regimes in the phase diagram—i.e., the spin liquid as well as the disordered state—which cannot be distinguished by a local order parameter due to the topological nature of the former.
@@ IV Discussion and outlook
In this work, we combine randomized measurements with tensor networks to perform quantum state tomography via MLE, with optional regularization using classical shadows. While the randomized measurements toolbox is useful for estimating many physical properties, tensor networks are known to efficiently represent physically relevant quantum states. Here, we have considered a more restricted measurement setting, random-$XZ$ shadows, that is well motivated by both its tomographic completeness on real and pure states and its feasibility in near-term experiments. Using these $XZ$ measurements, we can regularize our MLE loss function with observables estimated via the classical shadow framework. We showcase the performance of our protocol by learning the topological ground states of the surface code Hamiltonian as well as the Rydberg Hamiltonian on the ruby lattice. For both of these states, our protocol with random-$XZ$ measurements can accurately reconstruct the target state up to a fidelity of over $0.95$ with under $10^{5}$ samples. In addition, we test our shadow regularization technique on these two applications: we find that requiring the stabilizer expectations values to agree with estimations from classical shadows further improves the fidelity for the surface code, whereas for the ruby-lattice Rydberg spin liquid, this only reduces the number of outliers observed during the optimization process. Additionally, although our current work focuses on perfect measurements, it will be important for future studies to consider measurement errors and carefully analyze their effects in our protocol, as the errors will—in practice—limit the best possible fidelity. While we do not systematically analyze such effects in this work, in Appendix H, we summarize in more details the common sources of measurement errors in Rydberg atom arrays.
@@
Using higher-dimensional TNs combined with randomized measurements would, in principle, allow us to learn any states of interest. In the present work, we have numerically benchmarked our protocol using MPS as our model prior, given its advantages of exact contractions and sampling. An immediate extension would be to generalize this model’s architecture to represent 22D area-law entangled states, such as string bond states [57] or projected entangled pair states (PEPS) [58]. Another important generalization of our work, beyond pure states, is to adopting matrix product operator (MPO) ansätze, specifically, the locally purified density operator (LPDO) approach developed for learning quantum processes [29]. We leave the numerical demonstrations for such tensor-network states to future work.
@@
It constitutes an interesting perspective to extend these methods to the study of fermionic quantum states, which are the focus of attention in the study of correlated electron systems. Our analysis could be extended to quantum gas microscope measurements of optical lattice Fermi-Hubbard systems [59] which yield snapshots with the charge and spin on each site. While in this setting, quench dynamics generated by noninteracting Hamiltonians are presumably most conceivable [60], it is possible that the spin degrees of freedom can be measured using a similar unitary ensemble such as our random-$XZ$ measurements. The variational optimization can be carried out using fermionic neural networks [61] or auxiliary wavefunctions [62, 63], replacing the MPS we have employed in the present work. To enhance the optimization process using neural-network ansätze, it would be especially interesting to extend recent results [42] on locally randomized single-qubit measurements to such fermionic systems.
@@
While our method can be generalized to quantum states realized in a wide class of quantum processors, more specifically, for Rydberg quantum simulators, it can be easily extended to learning logically encoded quantum states. In particular, our random-$XZ$ measurements are well suited for near-term encodings, such as the color codes recently implemented with neutral atom arrays [31]. In such a Rydberg array, every logical qubit, encoded, for example, in a Steane code, can be measured in either the $X$ or the $Z$ basis, allowing for both error detection via post-selection and the ability to perform high-fidelity random-$XZ$ measurements. Even if current experiments do not aim to perform full logical quantum state tomography, our characterization and derivation of the tools required for random-$XZ$ classical shadows can be immediately applied to these logical encodings.
@@ V Code and data availability
The numerical code used for producing results and generating the figures in this work is available at: https://github.com/teng10/tn-shadow-qst.
@@ Acknowledgements.
We thank Jacob Barandes, Christian Bertoni, Jonathan Conrad, Zoë Holmes, Robert Huang, Dmitrii Kochkov, Nathan Leitao, Alexander Nietner, Jakob Unfried, Guifre Vidal, and Manuel Rudolph for useful discussions. We thank Hong-Ye Hu and Leandro Aolita for feedback on the manuscript. Y.T. is grateful to Dmitrii Kochkov for his inputs on the initial prototype of the numerical code base and systematic code reviews. Y.T. also thanks the groups of Zoë Holmes and Jens Eisert for their hospitality during her visits, when part of this work was completed. We thank the anonymous referee for pointing out the subtlety of measurement noise for states prepared via analog simulation. R.S. is supported by the Princeton Quantum Initiative Fellowship. K.V.K. acknowledges support from the Fannie and John Hertz Foundation and the National Defense Science and Engineering Graduate (NDSEG) Fellowship. The Berlin team acknowledges support from the BMBF (MuniQCAtoms), the DFG (CRC 183, FOR 2724), the Munich Quantum Valley, and the ERC (DebuQC). Y.T. and S.S. have been supported by the U.S. Department of Energy under Grant DE-SC0019030.
@@
This appendix aims to provide a comparison of our method to existing quantum state tomography techniques. We highlight the unique challenges and opportunities in characterizing complex quantum systems, particularly two-dimensional topological states. In addition to the references summarized below, we direct interested readers to Ref. [66] for a comprehensive review of quantum state tomography. Finally, having discussed the advantages of our method for learning quasi-22D quantum states with topological order, we summarize the limitations faced by our method discussed in the main text.
@@
    Matrix product state (MPS) tomography [18, 19, 67, 20] is an effective method for reconstructing injective one-dimensional quantum states with short-range correlations associated with a local parent Hamiltonian, by finding the best MPS consistent with local reduced density matrices from randomized measurements. In principle, these reduced density matrices can be estimated from Pauli measurements using the techniques of classical shadows [68]. Its application to 22D topological states, however, is challenging: first of all, the size of the reduced density matrices required for uniquely determining a quantum state in 22D is unclear, and secondly, due to nontrivial ground-state degeneracies of topological states, they cannot be uniquely determined by local operators and thus, cannot be efficiently determined with this method.
@@
    Variational MPS tomography [26, 27, 28] extends MPS tomography to one-dimensional states without the injectivity constraint. It uses a variational approach to optimize the MPS representation through maximum likelihood estimation (MLE). Prior work also considered using matrix product operators (MPOs) for quantum process tomography [29]. Our work extends this protocol to 22D states using a “snake path”, commonly adopted in cylindrical DMRG calculations for 22D systems. We further enhance our method by incorporating regularization techniques inspired by classical shadow tomography, combining the benefits of both approaches to better characterize two-dimensional quantum systems. We refer the reader to Sec. II for the details of our protocol.
@@
    Neural network quantum state tomography (NNQST) [23] utilizes neural networks as variational ansätze within the MLE framework to model quantum states. Although particularly successful for states with fully positive amplitudes (for example, ground states of frustration-free stoquastic Hamiltonians [32]), this method requires careful consideration when the target states have nontrivial signs and phases. Using measurement outcomes from different bases is therefore essential—a requirement directly addressed in MPS-based approaches, which can inherently relate measurements post local unitary transformations to learn relative phases. In this context, it may be useful to explore a recently proposed architecture that can accurately represent quantum spin liquids by incorporating physical symmetries [69].
@@
    Generative modeling for density matrix reconstruction [24] extends the use of neural networks to mixed-state tomography. This approach uses single-qubit positive operator-valued measure (POVM) measurements for information-complete reconstructions. While generative models can theoretically model measurement probability distributions across an exponentially large outcome space, practical challenges arise in ensuring convergence to the accurate density matrix, particularly given the sparse nature of such distributions in large systems. It is unclear how effective this approach is [25] for learning highly entangled quantum states such as the topological ordered states considered in this work.
@@ A.1 Potential advantages and limitations
Our variational tensor network tomography, aimed at diagnostics and recovery of quantum states, offers potential for extracting various properties such as the von Neumann entanglement entropy, difficult to learn using other techniques. This method combines maximum likelihood estimation of the quantum state with randomized measurements and is detailed in previous works such as [26, 28, 29]. We demonstrate this approach using simplified $XZ$ measurements, which proves sufficient for the two-dimensional topological quantum states considered, with complex long-range entanglement. Such quantum states are represented by noninjective MPS, and can not be learned via MPS tomography [18, 19].
One significant limitation is that the target state must admit an efficient tensor-network representation. In numerical experiments, we use an MPS ansatz, recognizing that while it effectively approximates one-dimensional states, it may not capture the full complexity of two-dimensional systems. This works sets the stage for future exploration of more sophisticated tensor networks. We also discuss the adaptation of statistical properties of MLE for quantum tomography, outlining both its potential and the challenges posed by the gauge freedom in MPS representations.
Another limitation is that real laboratory states can be mixed, which requires using generalizations of tensor networks such as locally purified matrix product operators [79]. Additionally, one generally does not have access to the ground-truth fidelity with the lab state. This necessitates methods like cross-validation as discussed in Appendix G or efficient fidelity certifications [42] to validate the trained model against empirical data.
@@
Our variational algorithm in the framework of maximum likelihood estimation (MLE) aims to find the MPS representation $|\psi({A})\rangle$ that is most consistent with measurements of the target state vector $|\phi\rangle$. Intuitively, we expect that the accuracy of the optimal MLE solution improves with more samples. The goal of this appendix is to analyze the accuracy of such a solution as a function of the number of samples $N$. More precisely, we provide a probabilistic upper bound of the infidelity 1F1-F by deriving a concentration inequality using the statistical properties of MLE. First, we summarize our main finding of the probabilistic bound in the theorem below:
@@
Now, a parameter shifted by a multiples of $2\pi$ still represents the same distribution $\mathcal{P}_{\theta_{0}}=\mathcal{P}_{\theta_{0}+2\pi}$. Therefore, $\hat{\theta}^{(N)}$ need not converge to $\theta_{0}$. In physics, the phenomenon where multiple parameters describe the same function is known as gauge redundancy. In this simple scenario, we can explicitly account for the periodicity in the optimal estimator, or equivalently impose “gauge fixing” conditions to select a unique representative, by restricting the parameter domain to $0,2\pi)$. However, for distributions parametrized by MPS, it is challenging in practice to fix the gauge degrees of freedom consistently. Specifically, transforming MPS to canonical forms using singular value decomposition only partially fixes the gauge. Additional unitary-gauge degrees of freedom remain in such a decomposition. Moreover, degeneracies in the singular values introduce redundancy within the subspace spanned by degenerate eigenvectors.